### v4

# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

Abstract
We report the discovery of a universal geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By analyzing the dissociation of localized wavepackets, we demonstrate that the $4\pi r^{2}$ radial volume element acts as a topological filter that inherently masks a significant portion of a system’s energy budget, imposing a fundamental peak-to-mean bound of $R_{E}<0.5$. We introduce an invariant scaling law, $\alpha=MQ/\zeta$, and prove that the resulting energy detection ratio is scale-invariant across twelve orders of magnitude, bridging attosecond molecular science and nuclear physics. We identify a universal “Geometric Zero-Point” at $R_{E}\approx 0.33$, which precisely replicates the 7 eV discrepancy in $H_{2}^{+}$ fragmentation. Furthermore, we show that the population of excited-state manifolds and the increase in nuclear localization ($\zeta$) provide a definitive geometric mechanism for the spectral broadening observed across atomic and subatomic scales. Remarkably, the spectral morphology derived from our scaling law aligns with the universal 1/3 energy landmark of historical beta decay, while the high-mass limit naturally accounts for the sharpening of alpha spectra. Our results suggest that “missing energy” is often a topological artifact of 3D geometry rather than an exclusive signature of undetected particles. This work establishes a universal master curve for energy reconstruction and identifies a “detection crisis” in highly localized systems, where the true interaction energy becomes effectively invisible to peak-centric calorimetry.
pacs:
32.80.-t,32.80.Rm,32.80.Fb
I Introduction
A central challenge in experimental physics is the reconstruction of the total energy budget from a detected energy distribution. Across disparate scales—from the electron-volts of molecular dissociation to the mega-electron-volts of nuclear transitions—a persistent anomaly remains: the primary spectral peak often captures only a fraction of the available energy. Historically, this “missing energy” has been the catalyst for some of the most profound discoveries in physics, most notably the hypothesis of the neutrino to resolve the energy deficit in beta decay Ellis and Wooster (1927).
As illustrated in our conceptual framework (Fig. 1), we propose that these discrepancies are linked by a universal topological mechanism. The Geometric Masking effect (Fig. 1, Row 1) arises from the inherent $4\pi r^{2}$ volume weighting of 3D space, which suppresses detection probability near the coordinate origin where potential energy is maximal. This results in a systematic shift of the detected kinetic energy release (KER) peak away from the true system mean.
In this work, we demonstrate that these anomalies share a common geometric origin defined by an invariant scaling law, $\alpha=MQ/\zeta$ (Fig. 1, Row 2). We prove that for any localized 3D wavefunction, the interplay between mass ($M$), localization ($\zeta$), and repulsive charge ($Q$) constrains the energy detection ratio. We identify a 0.33 geometric landmark—a universal attractor where the detected peak represents approximately one-third of the total energy budget. This scaling law predicts that as particles become more localized, the detected energy decays precipitously, creating a “detection crisis” where the true interaction energy becomes effectively invisible to peak-centric calorimetry.
The predictive power of this framework is underscored by the striking alignment between our scaling law and diverse experimental anomalies (Fig. 1, Row 3). In attosecond calorimetry, our model accurately predicts the observed 7 eV “missing energy” in $H_{2}^{+}$ fragmentation, yielding a detection ratio of $R_{E}\approx 0.32$ that matches modern experimental data Zhu (2020, 2021) and accounts for the spectral broadening induced by excited-state populations. This molecular landmark is mirrored almost perfectly in the subatomic regime; our $P(E)$ distribution provides a robust fit for the historical 210Bi beta decay spectrum within the non-relativistic limit, where the average electron energy remains tethered to approximately one-third of the endpoint energy Ellis and Wooster (1927). The persistence of this 0.33 ratio across twelve orders of magnitude, regardless of the specific interaction or mass scale, identifies the geometric volume element as the universal arbiter of energy detection. Ultimately, this framework provides a unified baseline for energy reconstruction, offering a robust geometric filter to distinguish between the intrinsic masking effects of 3D quantum geometry and the true signatures of new physics.
Figure 1: Universal Invariant Scaling and the Geometric Masking of Kinetic Energy. (Row 1) Physical mechanism: The transition from a localized bound state (Z-Z) to a repulsive continuum (+Q+Q) is filtered by the $4\pi r^{2}$ radial volume element, shifting the detected energy peak ($R_{E}$) significantly below the true system budget. (Row 2) Mathematical framework: The invariant scaling law defines a universal manifold where energy detection ratios depend solely on the coupling of localization ($\zeta$) and repulsion ($Q$). (Row 3) Empirical validation: The 1/3 geometric landmark ($R_{E}=0.33$) provides a unified explanation for the 7 eV "missing energy" in attosecond $H_{2}^{+}$ fragmentation and the historical average-to-endpoint discrepancy in 210Bi beta decay.
II The Geometric Physics of Fragmentation
The core of our model rests on the sudden approximation, which remains valid when a system’s Hamiltonian changes on a timescale $\tau_{perturb}\ll\hbar/\Delta E$. Under this condition, the initial wavefunction $\psi_{0}(\mathbf{r})$ remains “frozen” at $t=0$, as there is insufficient time for the probability density to reorganize Makarov (2016). In the context of our investigation, we model the fragmentation as an instantaneous transition from an attractive potential to a repulsive Coulomb configuration:
$V_{initial}(r)=-\frac{Z_{att}}{r}\longrightarrow V_{final}(r)=+\frac{Q}{r}.$ (1)
This approximation captures the essential dynamics of ultrafast processes across scales, from attosecond molecular dissociation to the loss of confinement in nuclear physics. To bridge the gap between initial state geometry and final energy observables, we utilize the de Broglie-Bohm formulation. The total energy $E(\mathbf{r})$ of a fragment at a specific coordinate $\mathbf{r}$ is the sum of the repulsive potential and the local kinetic energy density $T_{local}(\mathbf{r})$. For a Slater-type orbital $R(r)=Nr^{n-1}e^{-\zeta r}$, the analytical expression for the conserved total energy is:
$E(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (2)
The observable energy distribution $P(E)$ is constructed by sampling initial positions according to the 3D radial probability density $\rho(r)=4\pi r^{2}|\psi_{0}(r)|^{2}$. Crucially, this $r^{2}$ volume weighting acts as a topological filter. In highly localized systems (high $\zeta$), the majority of the probability density is concentrated at small $r$, where the local kinetic energy is highly negative. This results in an energy distribution where the most probable detected energy ($E_{peak}$) is shifted significantly lower than the mean system energy.
III Numerical Evolution and Parametric Sensitivity
To evaluate the predictive power of our 3D Bohmian-based model, we calculated the resulting energy distributions $P(E)$ and compared them against full quantum simulations. While detailed comparisons with tRecX Scrinzi (2022) Time Dependent Schrödinger Equation (TDSE) benchmarks are provided in Appendix A.2, we note that our analytical local-energy approach precisely captures the primary experimental signatures—specifically the peak energy $E_{\text{peak}}$—across varying regimes. Notably, the consistency in peak prediction (see Fig. 5(a)) confirms that our simplified method effectively captures the essential physics of the 3D Coulomb explosion.
III.1 Parametric Sensitivity and Distribution Morphology
Figure 2: Parametric evolution and universal scaling of the peak-to-mean energy ratio. (a) Sensitivity of the energy distribution $P(E)$ to repulsive Coulomb charge $Q\in[0,2]$ for a fixed $1s$ orbital ($\zeta=1,l=0$). (b) Evolution of $P(E)$ with orbital exponent $\zeta\in[1,4]$ at fixed $Q=1$. (c) Evolution of the energy ratio $R_{E}=E_{\text{peak}}/\langle E\rangle$ as a function of $Q$ (at $\zeta=1$, black curve) and $\zeta$ (at $Q=1$, red curve), demonstrating convergence toward $0.5$ in the repulsion-dominated regime and an efficiency drop in the localization-dominated regime. (d) Two-dimensional contour map of $R_{E}$ across the $(\zeta,Q)$ parameter space. The color scale indicates the magnitude of the detection discrepancy, strictly bounded by the 3D geometric masking limit.
To characterize the influence of physical parameters on the detected energy, we systematically varied the repulsive Coulomb charge $Q$ and the orbital exponent $\zeta$ (see Fig. 2). As illustrated in Fig. 2(a), increasing the repulsive charge $Q$ leads to a linear shift in the distribution $P(E)$ toward higher energies. This is accompanied by a moderate broadening of the spectral width, as a larger $Q$ increases the potential gradient across the spatial extent of the orbital.
In Fig. 2(b), we examine the effect of the orbital exponent $\zeta$. As $\zeta$ increases, the wavefunction becomes more tightly bound, simultaneously increasing the local kinetic energy density ($T_{\text{local}}\propto\zeta^{2}$) and placing the particle density in regions of higher potential energy. Interestingly, the distribution broadens much more rapidly with respect to $\zeta$ than with $Q$. This suggests that the internal “quantum pressure” of the initial state plays a more dominant role in determining the final energy dispersion than the external repulsive force itself. The shape of these curves confirms that fragment detectors fundamentally probe the radial geometry of the initial state rather than the integrated energy mean.
III.2 The Energy Detection Ratio $R_{E}$ and the Detection Crisis
To quantify the discrepancy between experimentally observed fragment signals and theoretical energy conservation, we define the dimensionless energy detection ratio:
$R_{E}=\frac{E_{\text{peak}}}{\langle E\rangle}$ (3)
where $E_{\text{peak}}$ is the mode of the distribution $P(E)$ and $\langle E\rangle$ is the integrated expectation value of the total energy. This ratio serves as a metric for the efficiency with which a detector captures the system’s total energy budget. Fig. 2(c) and 2(d) illustrates the systematic evolution of $R_{E}$ across the parametric space.
III.2.1 High-Charge Convergence and Classicalization
As shown by the black curve in Fig. 2(c), $R_{E}$ increases monotonically with the repulsive charge $Q$, asymptotically approaching a limit of $0.5$. In the high-$Q$ regime, the repulsive potential $V=Q/r$ dominates the local kinetic energy terms, “classicalizing” the landscape and shifting the peak of the distribution closer to the statistical mean. However, the fact that $R_{E}$ remains strictly bounded below $0.5$ underscores that the 3D quantum spatial spread is an inescapable constraint on the fragment energy spectrum.
III.2.2 The Shell-Dependent Detection Crisis
A precipitous decline in $R_{E}$ is observed as the orbital exponent $\zeta$ increases (red curve in Fig. 2(c)). In our framework, $\zeta$ serves as a proxy for the effective nuclear charge and localization; small $\zeta$ values represent diffuse valence orbitals, while large $\zeta$ characterize highly localized core orbitals. As $\zeta$ increases, the quadratic scaling of the “quantum pressure” near the nucleus ($T_{\text{local}}\propto\zeta^{2}$) drives the integrated mean energy $\langle E\rangle$ to values much higher than the peak $E_{\text{peak}}$, which is primarily determined by the most probable radial position.
Our findings reveal a “shell-dependent detection crisis.” For valence-shell Coulomb explosions (low $\zeta$), $R_{E}$ is relatively high, indicating that the detected peak provides a reasonably transparent window into the system’s potential energy. Conversely, inner-shell events (high $\zeta$) are subject to a massive geometric shift; for $\zeta\approx 5$, the detector may observe as little as 15% of the total energy mean. Given that localization typically scales with particle mass, the extreme confinement of nucleons in nuclear physics implies a significantly suppressed $R_{E}$. The global contour map in Fig. 2(d) synthesizes these trends, confirming that the deeper a particle is “buried” in the atomic potential, the more the 3D quantum geometry masks the true interaction energy from classical detection.
IV The Dimensionless Master Curve and Scaling Symmetry
Figure 3: Universal Scaling and the 0.33 Geometric Landmark. Invariant scaling and the universal 0.33 geometric landmark. (a) Evolution of the peak-to-mean ratio $R_{E}$ as a function of repulsive charge $Q$ under the constraint $\zeta=Q,M=1,n=1$. The dashed horizontal line indicates the 0.33 stability limit. (b) $R_{E}$ as a function of orbital exponent $\zeta$ for the case $Q=1,M=\zeta,n=1$, exhibiting identical convergence to the 0.33 landmark. (c) Three-dimensional contour representation of the $(\zeta,Q)$ parameter space for $(\zeta,Q)\in[1,5]$. The visualization shows three invariant planes corresponding to dimensionless scaling factors $\alpha=0.5$ ($R_{E}\approx 0.26$), $\alpha=1.0$ ($R_{E}\approx 0.33$), and $\alpha=2.0$ ($R_{E}\approx 0.39$), calculated with the mass scaling $M=\alpha\zeta/Q$. (d) Explicit mapping of the energy ratio $R_{E}$ as a function of the dimensionless coupling constant $\alpha$. The baseline computation uses parameters $M=\zeta=Q=n=1$.
The emergence of the 0.33 landmark is rooted in a fundamental scaling symmetry of three-dimensional fragmentation dynamics. To characterize this, we examine the peak-to-mean ratio $R_{E}$ across the parametric landscape of mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). As demonstrated in Fig. 3, the ratio is governed by a single dimensionless coupling constant $\alpha$ defined by:
$\alpha=\frac{MQ}{\zeta}.$ (4)
For a $1s$ Slater orbital, substituting the scaling relation for the dimensionless coupling reveals that the local kinetic density scales as $\rho_{T}\propto\zeta^{2}/M=\frac{1}{\alpha}\zeta Q$, while the local potential density follows $\rho_{V}\propto Q\zeta$ (see Sec. A.4). By transforming to a dimensionless radial grid $r_{s}=\zeta r$, both energy components at every coordinate point scale linearly with the product $\zeta Q$ for any fixed $\alpha$. This local synchronization ensures that the geometric morphology of the resulting $P(E)$ distribution is invariant to the absolute energy scale, depending solely on the coupling constant $\alpha$. This is confirmed in Fig. 3(a), where normalized distributions $P(E/Q\zeta)$ for varying $Q$ and $\zeta$ collapse onto identical master curves. As $\alpha\to\infty$, the distribution narrows toward a fixed peak position, providing a robust, scale-invariant baseline for extracting intrinsic parameters ($M,\zeta,Q$) from fragmented experimental spectra.
Furthermore, this symmetry demonstrates that the detection ratio $R_{E}$ remains invariant along $M\propto\zeta$ trajectories. The system undergoes a self-similar expansion in energy space where the $4\pi r^{2}$ masking effect is preserved as a constant geometric partition. The invariance observed in the parallel trajectories of Fig. 3(b) and the invariant planes of Fig. 3(c) is therefore a direct consequence of this homogeneous scaling. At the $\alpha\approx 1$ equilibrium, this collapse manifests as a horizontal 1D plateau at $R_{E}\approx 0.33$, identifying the landmark as a structural constant of 3D quantum geometry rather than a dynamical variable.
To determine the universal values of $R_{E}$ across the scaling parameter $\alpha$, we utilize this invariance to map the ratio onto our previously computed parametric results:
$R_{E}(\alpha)=R_{E}(Q)\rvert_{M=\zeta=1}.$ (5)
As shown in the master curve in Fig. 3(d), the numerical behavior of $R_{E}$ is characterized by the following limits:
$\lim_{\alpha\to 0}R_{E}=0,\quad\lim_{\alpha\to\infty}R_{E}=0.5.$ (6)
When $\alpha=1$, the system achieves a “geometric equilibrium” where the $4\pi r^{2}$ volume element sequesters exactly two-thirds of the total energy in the high-momentum radial tail, yielding the 0.33 landmark. As $\alpha$ increases (representing high $Q$ or low $M$), the ratio monotonically climbs toward the 0.5 universal bound, a regime where the repulsive field dominates the initial-state quantum pressure.
Conversely, as $\alpha$ increases (representing high localization $\zeta$), $R_{E}$ decays precipitously, falling below 0.2 for $\zeta>5$. This signals a profound “detection crisis” where the vast majority of interaction energy becomes experimentally invisible to peak-centric calorimetry. This confirms that for sufficiently localized states—such as core-shell electrons or compressed nucleons—the true interaction energy is effectively masked by the topological constraints of 3D space. Consequently, the mapping $R_{E}(\alpha)$ serves as a universal master curve for energy reconstruction across all scales of 3D quantum fragmentation.
V Discussion: From Molecular Calorimetry to Beta Decay
V.1 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy
Figure 4: Energy distribution analysis and empirical benchmarks. (a) Proton KER spectra for $H_{2}^{+}$ dissociative ionization. Numerical results represent 1D (dashed lines, Ref. Yue and Madsen (2013)) and 3D (solid lines, this work) computations. Experimental data are denoted by triangle markers (Ref. Pavičić et al. (2005)) and dot markers (Ref. Rudenko et al. (2005), after applying a logarithmic transformation). Black curves correspond to ground-state ($n=1$) contributions; red curves include the summation of populated excited states. (b) Energy distribution $P(E)$ as a function of principal quantum number $n$ ($n=1,2,3$) for fixed $Q=1,\zeta=1$. (c) Comparison of the historical 210Bi beta-decay spectrum (triangle markers, Ref. Ellis and Wooster (1927)) with the theoretical $P(E)$ distribution at $\alpha=0.16$ (black dashed line).
The physical validity of the 0.33 geometric landmark is most strikingly evidenced by the long-standing energy discrepancy in attosecond $H_{2}^{+}$ fragmentation. As shown in the Fig. 4(a), traditional 1D semiclassical models Yue and Madsen (2013) predict a KER peak between 10101313 eV, significantly overestimating experimental observations. In contrast, our 3D simulations (solid lines) yield a primary peak near $3$ eV, in excellent agreement with experimental calorimetry Pavičić et al. (2005); Rudenko et al. (2005).
Our model replicates the observed $\approx 7$ eV shift in $H_{2}^{+}$ fragmentation without requiring external field effects or multi-photon dynamics, identifying the $4\pi r^{2}$ volume element as the primary driver of the discrepancy. While the large mass of the escaping protons ($M_{p}\approx 1836$ a.u.) would typically drive the system toward a classical energy ratio, their initial localization within the $1\sigma_{g}$ molecular manifold provides a critical symmetry-induced compensation. Because the protons are initially constrained along the internuclear axis—a spatial distribution templated by the axial $\psi\propto r\cos^{m}\theta e^{-\zeta r}$ symmetry—the angular components of the Laplacian ($\nabla^{2}\psi\propto m/r^{2}$) significantly enhance the local "quantum pressure." This effectively neutralizes the mass scaling in our dimensionless coupling $\alpha=MQ/\zeta$. This mechanism maintains the heavy fragment system near the $\alpha\approx 1$ geometric equilibrium, where the volume weighting sequesters the majority of the potential energy. Consequently, the detected proton energy represents only one-third of the integrated mean ($\langle E\rangle\approx 11$ eV), confirming that the "missing energy" in molecular calorimetry is a direct manifestation of 3D quantum geometry.
This interplay between symmetry and mass is further illuminated when compared to subatomic regimes. While axial alignment in molecules provides a "quantum pressure" that counteracts fragment mass, nuclear fragmentation generally lacks this directional constraint. In the spherically symmetric environment of the nucleus, the large fragment mass becomes the dominant driver of the scaling constant $\alpha\to\infty$, shifting the detection ratio toward the sharp-peaked classical limit observed in Alpha decay.
V.2 Nodal Structures and the Mechanism of Spectral Broadening
Beyond the ground state ($n=1$), higher principal quantum numbers introduce additional radial nodes that fundamentally alter the topology of the energy distribution $P(E)$. As illustrated in the Upper Left panel of Fig. 4(b), simulations for $n=1,2,$ and $3$ (at fixed $Q=1,\zeta=1,M=1$) reveal that while the position of the primary low-energy peak remains relatively stable, increasing $n$ triggers the emergence of secondary features at higher energies. Physically, the local kinetic energy $T_{\text{local}}(r)$ incorporates terms related to the curvature of these additional radial nodes, redistributing probability density into regions of higher local kinetic energy and creating a multi-modal distribution. In the high-$n$ limit, this secondary feature can undergo a dominance shift, eventually surpassing the primary peak as the most prominent characteristic of the spectrum. We note that the varying high-energy truncation points in these numerical curves are artifacts of grid-spacing limits during the discretization of increasingly complex orbital geometries.
This nodal redistribution provides a definitive geometric explanation for the spectral broadening observed in high-intensity $H_{2}^{+}$ experiments in Fig. 4(a). When excited states ($n>1$) are populated, the 3D model identifies secondary high-energy features between 4455 eV (red solid curve), which are entirely absent in 1D approximations. These features serve as a direct signature of the radial nodal structures in the initial-state wavepackets. As shown in Fig. 4(a), the inclusion of these excited-state contributions explains the characteristic broadening and structured high-energy tails observed in experimental data (black and red dot markers, Ref. Rudenko et al. (2005)). Consequently, while the $1/3$ landmark defines the primary detection peak, the population of higher shells ($\zeta_{\text{eff}}$) provides the underlying geometric mechanism for spectral width evolution and the observed calorimetric structure of the fragmentation.
V.3 The Geometric Origin of Missing Energy in Subatomic Physics
The identification of the $M=\alpha\zeta/Q$ scaling law provides a provocative new lens for interpreting energy anomalies in nuclear physics. Historically, the continuous electron spectrum observed in beta decay—characterized by an average energy $\langle E\rangle\approx 1/3E_{\text{max}}$—was the primary catalyst for the neutrino hypothesis Ellis and Wooster (1927). Notably, the morphology of the beta spectrum (Fig. 4(c)) is remarkably well-replicated by our model in the low-to-mid energy range ($0.5E_{\text{max}}$) using a coupling constant of $\alpha=0.16$. This value corresponds to a high-localization ($\zeta$) regime, consistent with the high atomic number ($Z$).
The divergence of the fit at higher energies is expected, as this region approaches the relativistic regime where the non-relativistic Laplacian operator used in our current $P(E)$ derivation overestimates the energy drive. Furthermore, the nuclear Q-value imposes a hard phase-space cutoff at the endpoint energy that truncates the theoretical high-energy tail. Despite these specific high-energy corrections, our results demonstrate that the $R_{E}=0.33$ ratio represents a fundamental geometric baseline for energy detection in localized 3D systems. This universal constraint offers a unified explanation for perceived energy suppression across diverse physical scales, from molecular Coulomb explosions to the quenching factors observed in dark matter direct detection experiments Levin and others (1991); Aprile and others (XENON Collaboration) (2023)
V.4 Scaling Limits: Comparing Alpha and Beta Decay
The universal nature of the scaling law is further underscored by the contrasting spectral morphologies of Alpha Krane (1987) and Beta decay. While the low-mass regime of Beta decay (210Bi) resides near the $\alpha\approx 1$ landmark—where 3D geometric masking is maximal—the emission of heavy helium nuclei in Alpha decay represents the high-$\alpha$ limit of our framework. As the scaling constant $\alpha=MQ/\zeta$ increases due to the massive fragment $M$, our model predicts a geometric sharpening of $P(E)$ and a convergence of $R_{E}$ toward the classical limit. This explains why Alpha spectra appear as nearly mono-energetic lines with high detection efficiency, in contrast to the broad, “missing energy” distributions of Beta decay. While Alpha transitions involve complex $Q$-value dynamics beyond our current potential-inversion framework, the shift from masked to visible energy peaks confirms that the observability of subatomic transitions is fundamentally governed by the dimensionless scaling of the 3D volume element.
V.5 Unified Mechanism of Spectral Broadening and Orbital Contraction
The spectral distribution $P(E)$ is critically governed by the localization parameter $\zeta$. As illustrated in our parametric study (Fig. 2), an increase in $\zeta$ induces a rapid broadening of the high-energy tail. In high-$Z$ nuclei, the proximity of inner-shell electrons effectively scales the $\zeta$ parameter, leading to the spectral broadening observed in traditional nuclear studies Wilkinson (1990).
For superheavy nuclei ($Z>100$), the extreme localization (high $\zeta$) increases the probability of detecting ultra-high energy fragments. This phenomenon aligns with experimental observations of relativistic orbital contraction Henning (2012). While currently interpreted via Fermi theory Pershina (2016), our model suggests these relativistic shifts are consistent with the geometric pressure inherent to high-$\zeta$ localized states, providing a scale-invariant link between atomic and relativistic nuclear dynamics.
VI Conclusion
In this work, we have demonstrated that the persistent discrepancy between theoretical energy budgets and experimental peaks in quantum fragmentation is not a technical limitation, but a fundamental topological artifact of three-dimensional space. By synthesizing localized Slater-type orbitals with full-dimensional TDSE benchmarks, we have established that the $4\pi r^{2}$ volume element acts as an inescapable geometric filter. This mechanism inherently masks a significant portion of a system’s interaction energy, imposing a universal peak-to-mean bound of $R_{E}<0.5$.
The discovery of the invariant scaling law, $\alpha=(MQ)/\zeta$, provides a unified framework for energy reconstruction across twelve orders of magnitude. We have shown that the 7 eV "missing energy" in attosecond $H_{2}^{+}$ fragmentation is a direct manifestation of this geometric constraint, resolving the anomaly without the need for external field corrections. Most significantly, the alignment between this 0.33 landmark and the historical energy morphology of beta decay suggests that the 1/3 ratio is a universal baseline for 3D detection.
Our findings provide a transformative perspective on the interpretation of “missing energy” in subatomic physics. By mapping different decay modes onto our universal master curve, we show that the discrepancy in Beta decay and the high-resolution efficiency of Alpha decay are two limits of the same geometric scaling. While the neutrino remains a cornerstone of the Standard Model, our results reveal that the baseline for energy detection in localized states is inherently suppressed by the topology of the initial state. Because localization ($\zeta$) typically increases with particle mass ($M$), subatomic systems are naturally driven toward regimes where $R_{E}$ is highly sensitive to 3D masking. Consequently, geometric masking must be rigorously decoupled from signatures of invisible flux. This work shifts the focus of future calorimetry from the search for undetected mass to the calibration of the intrinsic geometric filters of our three-dimensional quantum world.
Appendix A Methods
A.1 Bohmian Local Energy Derivation
The local kinetic energy is derived from the Laplacian of the initial spatial wavefunction: $T_{local}(\mathbf{r})=-\frac{1}{2M}\frac{\nabla^{2}\psi_{0}(\mathbf{r})}{\psi_{0}(\mathbf{r})}$. In spherical coordinates, this yields the radial dependence utilized in our $P(E)$ construction. Because each Bohmian particle conserves its total energy in a time-independent repulsive field, the distribution at $t=0$ is identical to the asymptotic kinetic energy distribution at $t\to\infty$.
A.2 Numerical Validation against TDSE
To validate the predictive power of the 3D model, we compared $P(E)$ against full quantum simulations performed via the tRecX TDSE package Scrinzi (2022). Using a 1s hydrogenic orbital ($\zeta=1$ and $l=0$), we found that peak positions show high consistency across all repulsive charges $Q=0,0.5,1$ (see Fig. 5(a)). Deviations in the distribution tails are attributed to the numerical constraints of grid-based solvers at $r\to 0$, whereas our analytical approach maintains precision in the near-nucleus region.
Figure 5: Energy Distribution Analysis. (a) Comparison of normalized energy distributions $P(E)$ between the present model (solid) and tRecX TDSE simulations (dashed) for $Q=0,0.5$ ($\zeta=1$). (b) Analytical fitting for $\zeta=1,Q=1$. (c) Results for $\zeta=5.2,Q=1.6$. The r-fit (green) and Gamma fit (cyan) identify the radial nature of the distribution. Note that $E_{peak}$ remains significantly lower than $E_{scatter}$, consistent with the expected energy reduction for higher $\zeta$ values in nuclear repulsion scenarios.
To identify a universal analytical form for the fragment energy spectrum, we performed a regression analysis on the energy distribution $P(E)$ using several statistical models. Fig. 5 illustrates the results for a ground-state hydrogenic orbital ($\zeta=1,Q=1$, Fig. 5(b)) and a more localized, higher-charge configuration ($\zeta=5.2,Q=1.6$, Fig. 5(c)). We compared the computed data (blue) against three candidate distributions: a radial-type distribution (r-fit), a Gamma distribution, and a $\chi^{2}$ distribution. As shown in in the figure, the r-fit (green) and the Gamma distribution (cyan) produce the highest fidelity, with a residual error of less than 2%. The near-perfect overlap between these two functions confirms that the energy distribution essentially obeys a radial distribution law.
A.3 Symmetry-Dependent Local Kinetic Energy
To evaluate the influence of orbital alignment on the energy ratio, we consider a generalized prolate orbital often encountered in molecular ionization:
$\psi(r,\theta)=\mathcal{N}r\cos^{m}\theta\exp(-\zeta r)$ (7)
where $n$ represents the degree of axial alignment. The Laplacian in spherical coordinates, $\nabla^{2}=\frac{1}{r^{2}}\frac{\partial}{\partial r}(r^{2}\frac{\partial}{\partial r})+\frac{1}{r^{2}\sin\theta}\frac{\partial}{\partial\theta}(\sin\theta\frac{\partial}{\partial\theta})$, applied to this state yields:
$\frac{\nabla^{2}\psi}{\psi}\approx\left(\zeta^{2}-\frac{4\zeta}{r}+\frac{2(1-m)}{r^{2}}\right)$ (8)
The local kinetic energy $T_{\text{local}}(r)=-\frac{1}{2M}\frac{\nabla^{2}\psi}{\psi}$ therefore incorporates a centrifugal-like term proportional to the alignment $m$. In $H_{2}^{+}$, the high value of $n$ effectively counteracts the large mass $M$ in the denominator, preventing the collapse of the energy distribution into the classical regime. This confirms that the 0.33 landmark is a robust attractor for both light spherical systems and heavy aligned systems.
A.4 Derivation of the Universal Scaling Law
To establish the physical basis for the 0.33 landmark, we examine the local energy densities of a system described by a $1s$ Slater-type orbital:
$\psi(r)=\sqrt{\frac{\zeta^{3}}{\pi}}e^{-\zeta r}$ (9)
The detected kinetic energy release (KER) is determined by the local mapping of potential $V(r)$ and kinetic $T(r)$ energy densities onto the $P(E)$ distribution, weighted by the 3D volume element $dV=4\pi r^{2}dr$. By transforming to a dimensionless scaled radial grid $r_{s}=\zeta r$, we reveal the universal scaling properties of the fragmentation morphology.
1. Potential Energy Scaling
For a repulsive Coulombic interaction $V(r)=Q/r$, the local potential energy density $\rho_{V}(r)=\psi^{2}(r)V(r)4\pi r^{2}dr$ is expressed in the scaled grid as:
$\rho_{V}(r_{s})=4\zeta Q\cdot r_{s}e^{-2r_{s}}dr_{s}$ (10)
This confirms that at every point on the universal grid $r_{s}$, the potential energy density scales as $\rho_{V}(r_{s})\propto Q\zeta$.
2. Kinetic Energy Scaling
The radial kinetic energy density is derived from the Laplacian operator $\hat{T}=-\frac{1}{2M}\nabla^{2}$. Acting on the $1s$ orbital, the operator yields:
$\hat{T}\psi(r)=-\frac{1}{2M}\left(\frac{d^{2}}{dr^{2}}+\frac{2}{r}\frac{d}{dr}\right)\psi(r)=-\frac{\zeta^{2}}{2M}\left(1-\frac{2}{\zeta r}\right)\psi(r)$ (11)
Transforming the kinetic density $\rho_{T}(r)=\psi(r)\hat{T}\psi(r)4\pi r^{2}$ to the scaled grid $r_{s}$ yields:
$\rho_{T}(r_{s})=-\frac{2\zeta^{2}}{M}\left(1-\frac{2}{r_{s}}\right)r_{s}^{2}e^{-2r_{s}}dr_{s}$ (12)
In this coordinate system, the local kinetic density scales as $\rho_{T}(r_{s})\propto\zeta^{2}/M$. By introducing the dimensionless coupling constant $\alpha=MQ/\zeta$, we substitute $M=\alpha\zeta/Q$ to obtain:
$\frac{\zeta^{2}}{M}=\frac{\zeta^{2}}{\alpha\zeta/Q}=\frac{1}{\alpha}Q\zeta$ (13)
3. Homogeneous Invariance
At the geometric equilibrium defined by $\alpha=1$, both potential and kinetic energy densities at every scaled grid point $r_{s}$ scale identically with the product $Q\zeta$:
$\rho_{V}(r_{s})\propto(Q\zeta)\cdot f(r_{s}),\quad\rho_{T}(r_{s})\propto(Q\zeta)\cdot g(r_{s})$ (14)
Since the ratio $\rho_{V}(r_{s})/\rho_{T}(r_{s})$ is independent of the absolute values of $Q$ and $\zeta$, the morphology of the resulting $P(E)$ distribution is self-similar. This mathematical homogeneity ensures that the detection ratio $R_{E}=E_{\text{peak}}/\langle E\rangle$ remains locked at the 0.33 landmark, manifesting as the invariant horizontal plateaus observed in our parameter space analysis.
A.5 $H_{2}^{+}$ Simulation Protocols and Parameter Selection
To validate our geometric model, full quantum-dynamical simulations of $H_{2}^{+}$ were performed using the tRecX software package. These simulations solve the three-dimensional TDSE to extract high-fidelity KER spectra.
Two distinct ionization regimes were utilized to benchmark the influence of initial-state geometry:
  1. 1.
    Ground-State Manifold: Simulations using linearly polarized laser fields were employed to isolate the ground-state dissociation pathway.
  2. 2.
    Excited-State Populations: Circularly polarized fields were utilized to drive enhanced population transfer into higher-lying electronic states (n>1). This protocol aligns with both experimental and theoretical findings Zhang et al. (2023); Geng et al. (2022); Rudenko et al. (2005), which demonstrate that circular polarization significantly broadens the KER spectrum by populating excited-state manifolds.
Consistent with established experimental conditions Pavičić et al. (2005), the laser parameters were set to a central wavelength of 800 nm with a peak intensity of $8\times 10^{13}W/cm^{2}$. The resulting 3D flux was integrated to produce the KER distributions compared in the lower panel of Fig. 4.